A lot runs between two parallel streets. The length of the lot on one of the streets is 80-ft. The length of the lot on the other street is 100-ft. The area of the lot is 18,000 square feet. How far apart are the two streets? (The formula for the area of a trapezoid is A=12(b1+b2)h
.)

Answer by math_tutor2020(3826)



A = (1/2)*h*(b1 + b2)
18000 = (1/2)*h*(80+100)
18000 = (1/2)*h*180
18000 = 90h
h = 18000/90
h = 200 feet